Week of 060206

Open Actions from last week: Run PIC FTS traffic from production castor2. Castor team to take c2sc3 cluster for internal tests.

Chair: H.Renshall

On Call: Maarten

Monday:

Log: Nothing

New Actions: T.Whibley to verify DNS changes 14.00-16.00 are agreed/happening. ML to stop c2sc3 queues and switch ASGC to production service. When done the castor team will take c2sc3 hardware for tests. It will be reincarnated after chep, India, in the higher caste rank of c2sc4 having had good karma as c2sc3.

Discussion: Over the weekend Sara, IN2P3 and BNL stopped their disk-tape runs on schedule. Sara and BNL needed to recuperate resources for production. FZK and NDGF had also stopped leaving DESY, PIC and ASGC (disk to disk). DESY announced stoppage this morning and ASGC will carry on using production services. O.Barring announced they are migrating Alice to their production castor2 instance today.

Tuesday:

Log:

New Actions:

Discussion:

Wednesday

Log: LFC has a problem on LHCb R/O catalog - sophie to investiagte

Actions:

  • Miguel to stop DB on grid8 for elonex intervention DONE

Discussion:

  • Miguel would like to know progress with RLS migr. - to check with sophie

Thursday

Log: Nothing

Actions:

Discussion:

Friday

Log:

Actions:

Discussion:

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2006-02-09 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback